OpenStudy (anonymous):

help pleaseeee !!! simplify (U^-3)^6

OpenStudy (anonymous):

what is \[-3\times 6\]?

OpenStudy (anonymous):



OpenStudy (anonymous):

i want the answer without negative exponent

OpenStudy (anonymous):

u - 729 Raise 3 to the 6th power u - 36 Simplify u - 36 u - 729

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

\[-3\times 6=-18\]!

OpenStudy (anonymous):

\[u^{-18}=\frac{1}{u^{18}}\]